OpenStudy (anonymous):

find dy/dx of y=ln(X/1+x^2)

OpenStudy (anonymous):

the derivative of ln(u)=\[\frac{ 1 }{ u }*\frac{ d }{ du }(u)\]

OpenStudy (anonymous):

can you use this rule?

OpenStudy (anonymous):

Yeah im just having trouble finding the u

OpenStudy (anonymous):

in your problem u=\[\frac{ x }{ 1+x ^{2} }\]

OpenStudy (anonymous):

then the ans will be \[\frac{ 1 }{ \frac{ x }{ 1+x ^{2} } }*\frac{ d }{ du }(\frac{ x }{ 1+x ^{2} })\]

OpenStudy (anonymous):

So I take the derivative of the u with a quotient rule?

OpenStudy (anonymous):

yes right

OpenStudy (raden):

or before you derive it, by using the property of log it can be y = ln(x/1+x^2) = ln(x) - ln(1+x^2) now you can derive it one by one